# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 21 Apr 2026 v4

@@ 5.1 The distribution of $\varphi(p-1)/(p-1)$
as does the continuity criterion $\sum_{f(\ell)\neq 0}1/\ell=\infty$. The same conclusions hold for $g(n)=\varphi(n)/n$.
@@
Since $|f(\ell)|\leq\log 2<1$ for every prime $\ell$, the series $\sum_{|f(\ell)|>1}1/\ell$ is vacuous. The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$. Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion is satisfied. The case $g(n)=\varphi(n)/n$ follows from $\log g(\ell)=-f(\ell)$. ∎
Proposition 2 (Deshouillers–Hassani; special case of [9, Proposition 2.1 and equation (2.1)]).
@@
$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\textup{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
@@
Apply Deshouillers and Hassani [9, Proposition 2.1] to
@@
By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [9, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
By the continuity criterion in the same proposition, $G_{m}$ is continuous if and only if
@@
diverges. Here $g(\ell)=1-\frac{1}{\ell}\neq 1$ for every prime $\ell$, and
@@
by the divergence of the prime harmonic series in every reduced residue class (see, e.g., [29, Corollary 4.10]). Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
@@
We shall also use the following standard Siegel–Walfisz (see [29, Corollary 11.21] or [39, Theorem II.8.17]) estimate in a range of polylogarithmic moduli.
Theorem 3 (Siegel–Walfisz).
@@
Lemma 4 (Density of even totient ratios).
@@
Theorem 4 (Limiting distribution and full support over primes).
@@
Apply Proposition 2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
@@
so $G_{2}$ is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes. Writing $G:=G_{2}$ (denoted $F$ in [9, Theorem 1.1]), there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
@@
Let $\mu_{\varphi}$ denote the limiting measure with distribution function $G$. Fix an even integer $m\geq 2$, and write
@@
By Deshouillers and Hassani [9, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$. If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
By Lemma 4, the support contains a dense subset of $0,\tfrac{1}{2}]$. Since the support is closed and already lies inside $0,\tfrac{1}{2}]$, it follows that
$\operatorname{supp}(\mu_{\varphi})=[0,\tfrac{1}{2}].$
Now let $0\leq a<b\leq 1/2$. Choose any $x\in(a,b)$. Since $x\in\operatorname{supp}(\mu_{\varphi})$, every neighbourhood of $x$ has positive $\mu_{\varphi}$-measure. Taking $r>0$ so small that $(x-r,x+r)\subset(a,b)$, we obtain
$\mu_{\varphi}((a,b])\geq\mu_{\varphi}((x-r,x+r))>0.$
@@
The support question for $\varphi(n)/n$ over the integers remains open; see Schoenberg [35] and [9].
5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Lemma 5 (Transport lemma).
@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Fix $\alpha\geq 0$ and $0<\varepsilon<1$. For all sufficiently large primes, $1-\varepsilon<a(p)<1+\varepsilon$, so
@@
Theorem 5 (Transfer to the PRIM-LWE density).
@@
where $G$ is the distribution function from Theorem 4.
@@
$P(p)=1+O\!\left(\frac{1}{p}\right)\qquad(p\to\infty\text{ through the primes}).$
By Theorem 4, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
5.3 Pure singularity
Having established that $c(p)$ shares the limiting distribution of $\varphi(p-1)/(p-1)$, we now investigate the fine structure of this law. By the Jessen–Wintner theorem [24], any limiting distribution arising from an infinite convolution of discrete measures must be of pure type—purely discrete, purely singular, or absolutely continuous. Continuity was already verified in Lemma 3, so the remaining question is whether the law is absolutely continuous or purely singular. The latter turns out to hold: the values of $c(p)$ cluster on a fractal-like set of Lebesgue measure zero, as we prove by adapting a classical argument of Erdős to the shifted-prime setting.
Lemma 6 (Shifted-prime pure-type and continuity criterion).
Let $f$ be a real strongly additive arithmetical function satisfying the three convergence conditions of the shifted-prime Erdős–Wintner theorem, so that f(p1)f(p-1) possesses a limiting distribution $\mu_{f}$ over the primes. Then $\mu_{f}$ is of pure type: it is either purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure. Moreover, $\mu_{f}$ is continuous (i.e. has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
@@ 5.3 Pure singularity
The relevant shifted-prime Erdős–Wintner framework is due to Kátai [25, Theorem 2] and Hildebrand [23, Theorem 5] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem [4] (see also [39, Theorem II.8.34]): the only property of the shift used in these proofs is the Bombieri–Vinogradov estimate for primes in arithmetic progressions, which holds uniformly for any fixed nonzero $a$. Hildebrand [23, p. 211] notes explicitly that the same methods apply to $\{p+a\}$ for arbitrary nonzero $a$. For the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [38, formula (3), p. 262]. The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
By Tenenbaum [38, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
$\widehat{\mu_{f}}(\tau)\;=\;\prod_{\ell\textup{ prime}}\!\left(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\right).$
The shift is p1p-1, corresponding to a=1a=-1 in the notation of [38, formula (3)]. Since no prime divides 1-1, the product runs over all primes. For the shift a=1a=-1, the local factor at $\ell=2$ in the characteristic function reduces, when $f$ is strongly additive, to $1-\frac{1}{2-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(2)}}{2^{\nu}}=e^{i\tau f(2)}$, which is a deterministic translation by $f(2)$. This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law. For each odd prime $\ell$, the remaining local factor is the characteristic function of an atomic probability measure $\nu_{\ell}$ placing mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$. The product formula therefore represents the characteristic function of the infinite convolution $\delta_{f(2)}\ast\mathop{{\ast}}_{\ell\geq 3}\nu_{\ell}$. By Lévy’s uniqueness theorem (see, e.g., [39, Section III.2]), $\mu_{f}$ equals this infinite convolution. It therefore suffices to apply the Jessen–Wintner theorem (see also Tenenbaum [39, Theorem III.2.7(b)]) to the convolution over the odd primes. Each factor $\nu_{\ell}$ is supported on a countable (indeed two-element) set, so the Jessen–Wintner theorem gives that the resulting infinite convolution—and hence $\mu_{f}$ itself—is of pure type. The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of $\nu_{\ell}$. Since $\nu_{\ell}$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$, when $f(\ell)\neq 0$ these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$. It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution. Since a deterministic translation preserves pure type and continuity, the same conclusions hold for $\mu_{f}$. ∎
@@
Lemma 7 (Prime density for a prescribed smooth kernel).
@@
$C_{y}:=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}\;\asymp\;\frac{1}{\log y},$
we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
@@
For each odd prime $\ell\leq y$ with $\ell\mid m$, the condition $\ell\mid(p-1)$ forces $p\equiv 1\pmod{\ell}$. This selects one residue class out of $\ell-1$ reduced classes modulo $\ell$. For each odd prime $\ell\leq y$ with $\ell\nmid m$, the condition $\ell\nmid(p-1)$ excludes the class $p\equiv 1\pmod{\ell}$, leaving $\ell-2$ admissible classes out of $\ell-1$. The prime $\ell=2$ imposes no constraint since $2\mid(p-1)$ for every odd prime $p$ and $2\mid m$ by hypothesis. By the Chinese Remainder Theorem, these local conditions combine to specify a union of reduced residue classes modulo the primorial $M_{y}:=\prod_{\ell\leq y,\;\ell\text{ prime}}\ell$. The finitely many primes $p\leq y$ need not satisfy these residue-class conditions, but they contribute $O_{y}(1)$ to the count and do not affect the relative density. Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 3) gives the stated density.
Since each $m$ is even and squarefree, $\varphi(m)=\prod_{\begin{subarray}{c}\ell\mid m\\ \ell>2\end{subarray}}(\ell-1)$, so the first product in the expression for $\delta_{y}(m)$ equals $1/\varphi(m)$. The lower bound follows because each factor $(\ell-2)/(\ell-1)\in(0,1)$. To justify the asymptotic $C_{y}\asymp 1/\!\log y$, write
$\log C_{y}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$. By Mertens’ second theorem [29, Theorem 2.7(d)] the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$. ∎
Lemma 8 (Tail-mean estimate for f(p1)f(p-1)).
Let $f$ be as in (2), and for each prime $p$ and parameter $y\geq 3$ define
@@
Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 3),
@@
$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}\;\ll\;\frac{x}{L}\;+\;\log\log x.$
@@
Lemma 9 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
Let $f$ be the strongly additive function defined in (2). The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure. Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
@@
We adapt the integer-case argument of Erdős [14, pp. 722–725] (see also [39, Exercise 256] with solution in [40, Solution 2.4, pp. 147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 3) in place of the integer equidistribution input.
By the shifted-prime Erdős–Wintner theorem (Lemma 6), the limiting measure $\mu_{f}$ exists. Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous. By Lemma 6, $\mu_{f}$ is of pure type. Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous. We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.

### 06 May 2026 v5

@@ 5.1. The distribution of $\varphi(p-1)/(p-1)$
as does the continuity criterion $\sum_{f(\ell)\neq 0}1/\ell=\infty$. The same convergence conditions and continuity criterion hold for f-f, and the limiting distribution of $g(p-1)=\varphi(p-1)/(p-1)$ is the push-forward of the limiting distribution of f(p1)f(p-1) under $u\mapsto e^{-u}$.
@@
Since $|f(\ell)|\leq\log 2<1$ for every prime $\ell$, the series $\sum_{|f(\ell)|>1}1/\ell$ is vacuous. The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$. Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion is satisfied. Since $\log(\varphi(n)/n)=-f(n)$ is also strongly additive with the same absolute values at primes, all three series conditions and the continuity criterion transfer to f-f. The distributional conclusion for $g(n)=\varphi(n)/n$ then follows by applying the continuous map $u\mapsto e^{-u}$. ∎
Proposition 5.2 (Deshouillers–Hassani; special case of [7, Proposition 2.1 and equation (2.1)]).
@@
$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\Bigl\{p\leq x:p\textup{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\Bigr\}\\ =G_{m}(\alpha).$
@@
Apply Deshouillers and Hassani [7, Proposition 2.1] to
@@
Since $\log g(\ell)=-\log(\ell/(\ell-1))=-f(\ell)$, Lemma 5.1 (which verifies the convergence conditions for both $f$ and f-f) shows that the three convergence hypotheses of [7, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
By the Jessen–Wintner continuity criterion (applied in [7, Proposition 2.1, equation (2.1)]), $G_{m}$ is continuous if
@@
diverges. (The full criterion requires $\sum_{\ell\nmid m,\,g(\ell)\neq 1}1/\ell=\infty$, which holds a fortiori since $g(\ell)\neq 1$ for every prime.) Here $g(\ell)=1-\frac{1}{\ell}\neq 1$ for every prime $\ell$, and
@@
by the divergence of the prime harmonic series in every reduced residue class (see, e.g., [24, Corollary 4.10]). Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
@@
We shall also use the following standard Siegel–Walfisz (see [24, Corollary 11.21] or [33, Theorem II.8.17]) estimate in a range of polylogarithmic moduli.
Theorem 5.3 (Siegel–Walfisz).
@@
Lemma 5.4 (Density of even totient ratios).
@@
Theorem 5.5 (Limiting distribution and full support over primes).
@@
Apply Proposition 5.2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
@@
so $G_{2}$ is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes. Writing $G:=G_{2}$ (denoted $F$ in [7, Theorem 1.1]), there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
@@
Let $\mu_{G}$ denote the limiting measure with distribution function $G$. Fix an even integer $m\geq 2$, and write
@@
By Deshouillers and Hassani [7, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$. If $x_{m}\notin\operatorname{supp}(\mu_{G})$, then some open interval $I$ containing $x_{m}$ has $\mu_{G}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{G})$ for every even $m$.
By Lemma 5.4, the support contains a dense subset of $0,\tfrac{1}{2}]$. Since the support is closed and already lies inside $0,\tfrac{1}{2}]$, it follows that
$\operatorname{supp}(\mu_{G})=[0,\tfrac{1}{2}].$
Now let $0\leq a<b\leq 1/2$. Choose any $x\in(a,b)$. Since $x\in\operatorname{supp}(\mu_{G})$, every neighbourhood of $x$ has positive $\mu_{G}$-measure. Taking $r>0$ so small that $(x-r,x+r)\subset(a,b)$, we obtain
$\mu_{G}((a,b])\geq\mu_{G}((x-r,x+r))>0.$
@@
Remark 5.6 (Alternative support argument).
Full support also follows from the Bernoulli-convolution representation established in Section 6.1: the support of the infinite convolution $\mu_{f}$ of Lemma 6.1 is the closure of the set of partial sums $\log 2+\sum_{\ell\in S}\log(\ell/(\ell-1))$ over finite sets $S$ of odd primes, and the density argument of Lemma 5.4 shows this closure equals $\log 2,\infty)$. The proof above uses instead the non-differentiability result of Deshouillers–Hassani, which connects the support question to the local analytic behaviour of $G$.
For the classical distribution of $\varphi(n)/n$ over the integers, the topological support is $[0,1]$: the argument of Lemma 5.4, applied over all primes and without the forced factor $1/2$, gives density of $\{\varphi(n)/n:n\geq 1\}$ in $[0,1]$, and the Erdős–Wintner theorem [11] (see also Schoenberg [30]) identifies this closure with the support. The global modulus of continuity remains open; see Section 9.
5.2. From $\varphi(p-1)/(p-1)$ to $c(p)$
Lemma 5.7 (Perturbation lemma).
@@ 5.2. From $\varphi(p-1)/(p-1)$ to $c(p)$
For $\alpha<0$ both sides are zero since $X(p)\geq 0$ and $a(p)>0$, so the claim is immediate. We may therefore assume $\alpha\geq 0$.
Fix $0<\varepsilon<1$. For all sufficiently large primes, $1-\varepsilon<a(p)<1+\varepsilon$, so
@@
Theorem 5.8 (Transfer to the primitive-root determinant density).
@@
where $G$ is the distribution function from Theorem 5.5.
@@
$P(p)=1-O\!\left(\frac{1}{p}\right)\qquad(p\to\infty\text{ through the primes}).$
By Theorem 5.5, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5.7 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
5.3. Pure singularity
Having established that $c(p)$ shares the limiting distribution of $\varphi(p-1)/(p-1)$, we now investigate the fine structure of this law.
Note. The strictly stronger result $\dim_{H}(\mu_{G})=0$—that is, the limiting measure is carried by a Borel set of Hausdorff dimension zero—is established independently in Section 6.1 by a different method (a Borel–Cantelli covering argument exploiting the independence structure of the Bernoulli convolution). That result implies pure singularity as an immediate corollary. We include the Erdős-style argument for its independent methodological interest: it provides the first complete adaptation of Erdős’s classical technique [10] to the shifted-prime setting.
The shifted-prime Erdős–Wintner theorem (Theorem 5.9) gives the existence and characteristic function of the limit law for any additive function on $\{p-1\}$ satisfying three convergence conditions; for strongly additive functions it also yields that the law is of pure type and identifies a continuity criterion. Since Lemma 5.1 verifies the hypotheses for $f=\log(n/\varphi(n))$ and the continuity criterion holds, the remaining question is whether the law is absolutely continuous or purely singular. The latter turns out to hold: we prove pure singularity by adapting a classical argument of Erdős to the shifted-prime setting.
Theorem 5.9 (Shifted-prime Erdős–Wintner theorem; Hildebrand [16], Jessen–Wintner [18]).
Let $f$ be a real-valued additive arithmetical function. The following are equivalent:
  1. (i)
    The distribution functions
    $\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;f(p-1)\leq z\bigr\}$
    converge weakly, as $x\to\infty$, to a distribution function $G(z)$.
  2. (ii)
    The three series
    (5.2)  $\sum_{\begin{subarray}{c}\ell\textup{ prime}\\ |f(\ell)|>1\end{subarray}}\frac{1}{\ell},\qquad\sum_{\ell\textup{ prime}}\frac{f^{*}(\ell)^{2}}{\ell},\qquad\sum_{\ell\textup{ prime}}\frac{f^{*}(\ell)}{\ell}$
    all converge, where $f^{*}(n):=f(n)\mathbf{1}_{|f(n)|\leq 1}$.
When these conditions hold, the characteristic function of the limit law is
(5.3) $\widehat{G}(\tau)\;=\;\prod_{\ell\textup{ prime}}\biggl(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\biggr).$
Furthermore, if $f$ is strongly additive (i.e. $f(\ell^{\nu})=f(\ell)$ for every prime $\ell$ and every $\nu\geq 1$):
  1. (a)
    (Pure type.) The limit law is of pure type: purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure.
  2. (b)
    (Continuity criterion.) The limit law is continuous (i.e. has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
@@ 5.3. Pure singularity
[…]
Concretely, Hildebrand’s Theorem 5 [16] is stated for the distribution functions $\nu_{x}(p;\,f(p+1)\leq z):=(1/\pi(x))\#\{p\leq x:f(p+1)\leq z\}$. Replacing $p+1$ by p1p-1 throughout changes only the residue class in the Bombieri–Vinogradov input:
$\pi(x;q,-1)\;\longrightarrow\;\pi(x;q,1),$
for which the same level of distribution holds (see, e.g., [16]). The resulting theorem for $\{p-1\}$ is therefore identical in statement and proof, and no new estimates are required.
The explicit characteristic function (5.3) is recorded by Tenenbaum [32, formula (3)] for a general nonzero shift $a$, confirming the Euler-product form directly. For a=1a=-1 no prime divides $a$, so the product in (5.3) runs over all primes $\ell$ without restriction.
Pure type and continuity (strongly additive case). Assume now that $f$ is strongly additive; then $f(\ell^{\nu})=f(\ell)$ for all $\nu\geq 1$, so each local factor in (5.3) reduces to the characteristic function of an atomic probability measure supported on $\{0,f(\ell)\}$. (For $\ell=2$ the weight at 0 vanishes and the factor degenerates to the point mass $\delta_{f(2)}$; for odd $\ell$ it is a genuine two-point measure.) By Lévy’s uniqueness theorem (see, e.g., [33, Section III.2]), the limit law equals the infinite convolution of these local factors. The Jessen–Wintner theorem [18] (see also [33, Theorem III.2.7(b)]) states that every infinite convolution of discrete probability measures is of pure type, giving (a).
For the continuity criterion (b), since $f$ is strongly additive each local factor reduces to $\frac{\ell-2}{\ell-1}\,\delta_{0}+\frac{1}{\ell-1}\,\delta_{f(\ell)}$, a two-point measure. Let $\sigma_{\ell}$ denote its maximal atom. The Jessen–Wintner theorem gives: the convolution is continuous if and only if $\sum_{\ell}(1-\sigma_{\ell})=\infty$. When $f(\ell)\neq 0$, the atoms at 0 and at $f(\ell)$ are distinct, so $\sigma_{\ell}=\max\!\bigl(\frac{\ell-2}{\ell-1},\,\frac{1}{\ell-1}\bigr)=\frac{\ell-2}{\ell-1}$ for $\ell\geq 3$, whence $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$. When $f(\ell)=0$, the two-point measure collapses to $\delta_{0}$ and $1-\sigma_{\ell}=0$. Therefore $\sum_{\ell}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$. ∎
@@
Lemma 5.10 (Prime density for a prescribed smooth kernel).
@@
$Q(y):=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}\;\asymp\;\frac{1}{\log y},$
we have $\delta_{y}(m)\geq Q(y)/\varphi(m)$.
@@
For each odd prime $\ell\leq y$ with $\ell\mid m$, the condition $\ell\mid(p-1)$ forces $p\equiv 1\pmod{\ell}$. This selects one residue class out of $\ell-1$ reduced classes modulo $\ell$. For each odd prime $\ell\leq y$ with $\ell\nmid m$, the condition $\ell\nmid(p-1)$ excludes the class $p\equiv 1\pmod{\ell}$, leaving $\ell-2$ admissible classes out of $\ell-1$. The prime $\ell=2$ imposes no constraint since $2\mid(p-1)$ for every odd prime $p$ and $2\mid m$ by hypothesis. By the Chinese Remainder Theorem, these local conditions combine to specify a union of reduced residue classes modulo the primorial $M_{y}:=\prod_{\ell\leq y,\;\ell\text{ prime}}\ell$. The finitely many primes $p\leq y$ need not satisfy these residue-class conditions, but they contribute $O_{y}(1)$ to the count and do not affect the relative density. Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 5.3) gives the stated density.
Since each $m$ is even and squarefree, $\varphi(m)=\prod_{\begin{subarray}{c}\ell\mid m\\ \ell>2\end{subarray}}(\ell-1)$, so the first product in the expression for $\delta_{y}(m)$ equals $1/\varphi(m)$. The lower bound follows because each factor $(\ell-2)/(\ell-1)\in(0,1)$. To justify the asymptotic $Q(y)\asymp 1/\!\log y$, write
$\log Q(y)=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$. By Mertens’ second theorem [24, Theorem 2.7(d)] the sum over primes equals $\log\log y+O(1)$, so $Q(y)=e^{-\log\log y+O(1)}\asymp 1/\!\log y$. ∎
Lemma 5.11 (Tail-mean estimate for f(p1)f(p-1)).
Let $f$ be as in (5.1), and for each prime $p$ and parameter $y\geq 3$ define
@@
Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 5.3),
@@
$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}$
$\;\ll\;\frac{x}{L}\;+\;\log\log x.$
@@
Lemma 5.12 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
Let $f$ be the strongly additive function defined in (5.1). The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure. Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
@@
We adapt the argument of Erdős [10, §3] (see also [34, Exercise 2.4]) to the shifted-prime setting, replacing the integer equidistribution input with the Siegel–Walfisz theorem (Theorem 5.3). Throughout, $y$ is fixed while $x\to\infty$; the moduli arising from the construction divide $M_{y}:=\prod_{\ell\leq y}\ell$ and hence lie in the Siegel–Walfisz range. Only after taking $x\to\infty$ do we let $y\to\infty$.
By Lemma 5.1 and Theorem 5.9, $\mu_{f}$ exists, is continuous, and of pure type. It remains to rule out absolute continuity.